# Uncertainty Quantification of Surrogate Models using Conformal Prediction

### v2

@@ 4.1 Strengths
In physical modelling and engineering design, ML models are deployed as emulators to replace expensive simulation codes and control algorithms, so that they can help efficiently and inexpensively perform design optimisation and uncertainty quantification (i.e., evaluating input data uncertainty and computing risks of undesirable scenarios). For these applications, ideally, the surrogate is trained once for simulation data and then deployed for a particular application with few options for further data acquisition and fine-tuning. However, in these cases, it is important to equip predictions with credible uncertainty estimates as more physics-based surrogate models evolve from a research environment to a production suite. In such critical modelling tasks, the validity of uncertainty plays a crucial role in determining the efficacy of a model (Begoli et al., 2019). Conformal prediction is suitable for these settings as it provides statistical guarantees for the coverage established by the framework. Aside from validity in coverage, CP provides the added benefit of application across pre-trained models, without requiring any knowledge about the model architecture (also not needing any architectural modifications) and its training regime. Within this paper, we showed that the CP framework can be formalised and set to obtain statistically valid marginal coverage over high-dimensional spatio-temporal outputs, beating the curse of dimensionality with near-zero computational costs. It provides statistical guarantees over the coverage of the prediction datasets while being relatively inexpensive and trivial in its implementation. CP also provides a methodology for uncertainty quantification that allows further testing of the model’s usefulness in a particular prediction regime.
@@ 4.2 Limitations
Many AI applications in scientific domains employ machine learning models that are trained and deployed without considering specific application contexts. These ML workflows often have constraints on retraining capabilities or obtaining additional data from simulations. The following paragraphs will explore potential limitations of this approach and discuss various solutions proposed in the literature.
Marginal coverage. The main prohibitive assumption is the fact the coverage guarantees (1) only provide marginal coverage, i.e., coverage is only valid for predictions averaged over all labels. For surrogacy, conditional coverage is desirable
$\mathbb{P}(Y_{n+1}\in\mathbb{C}^{\alpha}|X_{n+1})\geq 1-\alpha,$ (6)
i.e., to wish for $\mathbb{C}^{\alpha}$ to be valid for each specific label the model predicts. An ideal framework would maintain coverage on each particular prediction, and not averaged over a dataset. Although this feature is not possible to guarantee in general, it can be approximated reasonably well (Vovk, 2012).
Within the formulation considered in this paper, it must be reiterated that the coverage is estimated and hence, guaranteed cell-wise across the predicted spatio-temporal tensor. It does not guarantee a joint coverage that extends across the full scale of the bounds of the interested prediction domain. Though there has been work done in extending from marginal coverage across each cell to joint coverage across each domain, they fail to scale with dimensionality (Diquigiovanni et al., 2021; Messoudi et al., 2021, 2022).
An alternative solution without re-evaluating a novel calibration dataset is to use covariate shifting (Tibshirani et al., 2019). An additional assumption is however required: both the probability density of the training calibration distribution and of the new distribution must be known or could be inferred effectively. This is a fairly weak assumption in surrogate modelling, where the training calibration distribution is certainly known (the distribution we originally trained our surrogate on), and the new distribution’s density could perhaps be estimated if only sample data is obtained. Probability density estimation required for performing CP under covariate shift fails to scale effectively with high dimensionality (Quionero-Candela et al., 2009).
Data Intensive. Though the framework does not require any retraining, it still requires additional simulation / experimental data to perform the calibration, which might be prohibitively expensive. The statistical coverage guarantees from the CP framework itself lie on a beta distribution as given in Equation 4, governed by the number of calibration data points. We further conduct an empirical study of the impact of calibration data on the obtained coverage which is provided in Appendix G.
No predictive distribution. Unlike Bayesian methods, which provide probability distribution as predictions, conformal prediction provides prediction sets. Distributions are simpler to compute with, as they may be simply evaluated or propagated by Monte Carlo sampling. They also provide precise probabilities for events or any derived risk calculation. Methods of performing calculations such as calibration, propagation, or computing probabilities of events with prediction sets are not obvious, which is perhaps a reason why frequentist methods are not as widely used in UQ for scientific applications.
Recently, however, Cella and Martin (2022) have proposed an imprecise probabilistic interpretation of conformal prediction, and describe how probability measures may be bounded in conformalised regression and classification problems. They describe how the contour function of a plausibility measure (i.e., a possibility measure) can be computed from the conformal prediction sets. This could allow for the uncertainty estimates from conformal prediction to be further propagated and evaluated (Balch, 2012; Hose and Hanss, 2021), and makes a wider range of uncertainty quantification tasks accessible to confidence-based methods like conformal prediction.
Although Bayesian posteriors are perhaps also more informative, Balch et al. (2019), they must be treated with caution for modelling epistemic uncertainty. Martin (2019) describes this in a mathematically rigorous way, through the false-confidence theorem, stating that any additive model of uncertainty (e.g., a precise probability) is not guaranteed to provide correct inferences on all hypotheses: that there are certain events that a posterior distribution will assign a high probability to when the true probability is low. Although predictive distributions are simpler to understand and work with, their use does not come without its challenges.
Input Independence. Uncertainty ($\sim\hat{q}$) is added to the model outputs as prediction sets are independent of the model inputs and a function of the calibration dataset. In more Bayesian methods, the prediction uncertainty, estimated in the posterior is seen as a function of the prior and the likelihood, hence varying for each region of the prediction regime. Within the deterministic models explored within this work (using nonconformity score AER), the value of the error ($\sim\hat{q}$) is determined before the model prediction and remains fixed across the inference regime. This is exceptionally clear in the case of the fine-tuned foundation physics model (see Figure 12, where the error bars indicate the global error approximation and not on a case-by-case basis. For probabilistic models (using STD nonconformity score) this is less severe as the error bar is not just a function of $\hat{q}$ but also of the standard deviation output from the model. This allows for the prediction sets to be more flexible and adjust as a function of the model inputs (see Figure 2). Though there has been work on making the CP framework further input dependent (Johansson et al., 2021), they still require augmenting deterministic models through ensembles or architectural changes.
@@ 5 Conclusion
In this paper, it is shown that by maintaining the exchangeability of the data across the spatio-temporal domain (i.e. by preserving the tensorial shape), conformal prediction provides valid error bars (i.e., satisfied Equation 1 to expensive surrogate models for scientific applications, even if they make out-of-distribution predictions. Our empirical studies demonstrate that conformal prediction can be applied to any pre-trained / fine-tuned model for obtaining statistically guaranteed coverage, irrespective of the model configuration, training conditions and dimensionality of the outputs. One of the crucial highlights of the study is that CP can help provide guaranteed prediction sets across a new set of PDE solutions (out-of-distribution data), different to that which the model was conditioned on. Our work benchmarks against three kinds of calibration methods (nonconformity scores), across deterministic and probabilistic models, ensuring validity for the coverage. Although we focus on providing guaranteed coverage for modelling spatio-temporal data, we believe that our findings extend to any models predicting outputs with fixes fixed tensorial output shapes and with exchangeable calibration and prediction regimes.
The code used for the study can be found in https://github.com/gitvicky/Spatio-Temporal-UQ
@@ 6 Acknowledgements
The authors would like to thank Anima Anandkumar and Zongyi Li from Caltech for their help with defining neural operators and extending them to complex physics cases. The authors would also like to thank Michael McCabe at the Flatiron Institute for his help in setting up the Multi-Physics Pretrained foundation physics model and extending it to a Fusion-relevant database. This research was supported by the Excellence Center at Linköping–Lund in Information Technology (ELLIIT). Our computations were enabled by the Berzelius resource at the National Supercomputer Centre, provided by the Knut and Alice Wallenberg Foundation.
@@ Appendix A Poisson Equation
The Poisson equation is solved with a finite difference scheme using the py-pde python package (Zwicker, 2020). Equation 7 is constructed as an initial-value problem, where a scalar uniform field is initialised across the domain and evolved until convergence. A dataset comprising different instances of the 1D Poisson equation is constructed by sampling for different initial values uniformly from within the domain: $u_{\text{init}}\in[0,4)$.
@@
Each MLP is trained to take in the scalar initial field along the 32-point spatial domain to output the final field at the steady state. For the case of STD, the architecture is modified with 1D dropout layers. Each model is trained for up to 1000 epochs using the Adam optimiser Kingma and Ba (2015) with a step-decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a quantile loss for the case of CQR and MSE loss in all other cases.
@@ B.1 Physics
As discussed in section 3.2, the dataset is built by solving the one-dimensional Convection Diffusion equation numerically. The physics of the equation, given by the various coefficients is sampled from a certain range as given in Table 2. Each datapoint, as in each simulation is generated with different Diffusion coefficients and wave velocities as described in section 3.2. Each simulation is run for 100 time iterations with a $\Delta t=0.0005$ across a spatial domain spanning [0,10], uniformly discretised into 200 spatial units in the x-axis. Once the simulations are run and the dataset is generated, we downsample the temporal discretisation from 100 to 20 by slicing across the domain. The sampling parameters governing the PDE solutions used for the training are given in Table 2 and that used for the calibration and prediction is given in Table 3.
@@
Diffusion Coefficient $(\alpha)$ $[\sin(\frac{x}{\pi}),\sin(\frac{x}{2\pi})]$ Continuous
@@ B.2 Model and Training
We train a U-Net to map the spatio-temporal evolution of the field variable taking in the first 20 time instances (T_in) to the next 10 time instances (T_out). For the case of the Convection-Diffusion Equation, we don’t deploy an auto-regressive structure but perform a mapping from the initial distribution to the later distribution. The U-Net architecture can be found in Table 4. For the case of STD, the architecture is modified with 1D dropout layers following each encoder and decoder of the U-Net. Though the values governing the evolution of Convection-Diffusion are relatively small, for better representation we normalise the value with a linear range scaling, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser Kingma and Ba (2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a quantile loss for the case of CQR and an MSE loss in all other cases.
@@
Diffusion Coefficient $(\alpha)$ $[\sin(\frac{x}{2\pi}),\sin(\frac{x}{4\pi})]$ Continuous
@@ C.1 Physics
The solution for the wave equation is obtained by deploying a spectral solver that uses a leapfrog method for time discretisation and a Chebyshev spectral method on tensor product grid for spatial discretisation (Gopakumar et al., 2023). The dataset is built by performing a Latin hypercube scan across the defined domain for the parameters $\alpha,\beta,\gamma$, which accounts for the amplitude and the location of the Gaussian peak, sampled differently for each simulation. We generate 2,500 simulation points, each one with its own initial condition and use 500 for training, 1,000 each for calibration and procuring the prediction sets. We train a 2D U-Net and an FNO to learn the evolution of wave dynamics.
The physics of the equation, given by the various coefficients is held constant across the dataset generation throughout as given in Equation 10. Each data point, as in each simulation is generated with a different initial condition as described above. The parameters of the initial conditions are sampled from within the domain as given in Table 5. Each simulation is run for 150-time iterations with a $\Delta t=0.00667$ across a spatial domain spanning $[-1,1]^{2}$, uniformly discretised into 33 spatial units in the x and y axes. Once the simulations are completed and the dataset is generated, we select the first 80 time instances of the evolution of each simulation to be used for training.

### v3

@@ 4 Discussion
We have demonstrated that conformal prediction provides a practical, theoretically grounded approach to uncertainty quantification for surrogate models across diverse spatio-temporal applications. Through comprehensive empirical evaluation spanning PDEs, fusion diagnostics, and weather forecasting, CP delivers statistically guaranteed marginal coverage regardless of model architecture, training regime, or output dimensionality. However, understanding both its capabilities and inherent limitations is essential for responsible application in scientific domains. Below, we discuss CP’s key strengths for practical deployment, followed by an honest assessment of its limitations and implications for real-world use cases
@@ 4.1 Strengths
In safety-critical applications such as fusion reactor design, climate modelling, and engineering optimisation, surrogate models must provide credible uncertainty estimates alongside their predictions (Begoli2019). Conformal prediction addresses this need by offering statistical guarantees for uncertainty quantification with several key advantages:
Statistical Guarantees. CP provides provable marginal coverage (Equation 1) regardless of model architecture, training regime, or output dimensionality. This validity is particularly crucial when surrogate models transition from research to production environments where retraining opportunities are limited.
Model-Agnostic and Scalable. The framework requires no architectural modifications or knowledge of training procedures, enabling application to pre-trained models. Our experiments demonstrate guaranteed coverage across outputs spanning up to 20 million dimensions (weather forecasting) with near-zero calibration costs, effectively circumventing the curse of dimensionality.
Computational Efficiency. Unlike ensemble methods or Bayesian approaches requiring extensive sampling, CP calibration is computationally trivial (Section 2.3). Calibration times range from seconds for low-dimensional problems to minutes for high-dimensional applications, performed on standard hardware without specialised computational resources.
Practical Utility. CP enables rigorous validation of a surrogate model’s usefulness for specific downstream applications, providing actionable uncertainty estimates for decision-making in risk-averse scenarios.
@@ 4.2 Limitations
While CP offers substantial benefits, several inherent limitations must be acknowledged for responsible application in scientific domains. We discuss these limitations, their practical implications, and potential mitigation strategies.
Marginal vs. Conditional Coverage. The coverage guarantee in Equation 1 provides marginal coverage—validity averaged over all predictions—rather than the more desirable conditional coverage:
$\mathbb{P}(Y_{n+1}\in\mathbb{C}^{\alpha}|X_{n+1})\geq 1-\alpha.$ (6)
In practice, this means that while 90% of predictions will be covered on average, any individual prediction may have substantially different actual coverage. Although conditional coverage cannot be guaranteed in general, approximations exist (vovk2012conditional).
Our formulation guarantees coverage cell-wise across the spatio-temporal tensor but does not provide joint coverage across the entire prediction domain. Extensions to joint coverage exist (diquigiovanni2021conformal; messoudiCopula2021; messoudiEllipsoidal2022) but fail to scale with dimensionality, limiting their applicability to the high-dimensional problems considered here.
Practical Impact: In risk-critical scenarios requiring reliable bounds for specific predictions (e.g., fusion disruption avoidance, extreme weather events), marginal coverage may be insufficient. Users must understand that individual predictions, particularly in distribution tails, may not achieve the target coverage. Our experiments on exchangeability violations (Section 3.8, Figure 21) demonstrate this sensitivity.
Mitigation: Covariate shifting methods (tibshirani2019conformal) can improve conditional coverage when probability densities of calibration and deployment distributions are known or estimable. However, density estimation becomes unreliable in high dimensions (dataset_shift_book). For critical applications, we recommend conservative $\alpha$ values and validation on held-out data similar to deployment conditions.
Data Requirements and Exchangeability. CP requires calibration data that is exchangeable with the prediction regime. The quality of coverage guarantees follows a beta distribution (Equation 4) governed by calibration set size. Our empirical study (Appendix G) shows that $n_{\text{cal}}\geq 1000$ typically provides reliable coverage.
Practical Impact: For experimental data (weather, fusion diagnostics), obtaining sufficient exchangeable calibration data can be challenging. The exchangeability assumption is particularly delicate for time-series data where distribution shifts are common. As demonstrated in Figure 21, violations lead to invalid coverage. Weather experiments (Section 3.7) assume minimal climate variation between calibration (2021) and prediction (2022) years—an assumption that may not hold for long-term climate change scenarios or extreme events outside the calibration distribution.
Mitigation: When calibration data is expensive (e.g., fusion experiments), fine-tuning scenarios offer a natural solution: training data doubles as calibration data since both represent the target distribution. For time-series applications, careful validation of exchangeability assumptions and sensitivity analyses (as performed for camera diagnostics) is essential.
Prediction Sets vs. Distributions. CP provides prediction sets (intervals) rather than full probability distributions. While Bayesian methods offer distributions that can be propagated via Monte Carlo sampling or used for risk calculations, CP intervals lack this flexibility.
Recent work (cella2022validity) provides imprecise probabilistic interpretations of CP, enabling uncertainty propagation (balch2012mathematical; hose2021universal). However, these methods are not yet widely adopted. It is worth noting that while Bayesian posteriors appear more informative, the false-confidence theorem (martin2019false) shows that precise probability models can assign high confidence to low-probability events.
Input Independence in Deterministic Models. For deterministic models using AER nonconformity scores, the error bars ($\sim\hat{q}$) are fixed during calibration and do not vary with inputs. As shown in Figure 12, this produces global error estimates rather than input-specific bounds. This limitation is less severe for probabilistic models (STD scores) where prediction sets scale with model-predicted standard deviations (Figure 2), providing a weak conditionining on the input.
Practical Impact: Deterministic CP may produce overly conservative bounds in easy regions and insufficient bounds in challenging regions. This is particularly evident in foundation models trained on diverse data, where local complexity varies substantially.
Mitigation: When possible, use probabilistic models or ensemble approaches that provide initial uncertainty estimates for STD-based CP. Alternatively, recent work on normalised CP (normalisedCP2021) offers more input-adaptive bounds, though at the cost of increased computational complexity or architectural modifications.
Spatial and Temporal Correlations. Our cell-wise calibration treats each spatio-temporal point independently, ignoring correlations between adjacent cells. This is a significant limitation for PDE-based systems where solutions exhibit strong spatial and temporal dependencies.
Practical Impact: While we achieve marginal coverage at each cell, the framework does not capture or leverage the correlation structure inherent in physical systems. This may lead to overly conservative joint coverage across regions or miss spatially coherent error patterns.
Mitigation: We implicitly rely on the surrogate model to learn spatial dependencies during training. Future extensions could incorporate spatial or temporal correlation structures into the CP framework, leading to joint coverage as we have explored in gopakumar2025calibrated.
4.3 Broader Context
Despite these limitations, CP offers a valuable and practical tool for UQ in scientific machine learning. The guaranteed marginal coverage, model-agnostic nature, and computational efficiency make it particularly suitable for validating pre-trained surrogate models before deployment. When combined with careful validation of assumptions (particularly exchangeability), CP provides actionable uncertainty quantification that can guide decision-making in safety-critical applications.
For optimal results, we recommend: (1) validating exchangeability assumptions through sensitivity analyses, (2) using conservative $\alpha$ values for risk-critical applications, (3) preferring probabilistic models or ensembles when input-dependent bounds are crucial, and (4) maintaining awareness of the marginal nature of coverage guarantees when interpreting individual predictions.
@@ 5 Conclusion
This paper presents a comprehensive empirical study demonstrating that conformal prediction provides statistically guaranteed uncertainty quantification for surrogate models across diverse scientific applications. By maintaining exchangeability of spatio-temporal data through preservation of tensorial structure, we achieve valid error bars satisfying Equation 1 for outputs spanning up to 20 million dimensions.
Key Contributions. Our work establishes that CP can be applied to any pre-trained or fine-tuned surrogate model—regardless of architecture (MLP, U-Net, FNO, ViT, GNN), training regime, or output dimensionality—to obtain guaranteed marginal coverage with near-zero computational cost. We benchmark three nonconformity scores (CQR, AER, STD) across both deterministic and probabilistic models, demonstrating consistent coverage across applications ranging from fundamental PDEs to operational weather forecasting and fusion diagnostics.
Critically, we show that CP provides valid prediction sets even for out-of-distribution scenarios where models are deployed on physics regimes different from their training distribution (wave equation at half-speed, Navier-Stokes at different viscosity, pre-trained foundation models on new physics). This capability is essential for validating surrogate model utility in production environments where retraining is infeasible.
Practical Impact. For scientific machine learning practitioners, our framework offers a rigorous method to assess whether a pre-trained model is suitable for a specific downstream application. The guaranteed coverage enables confident deployment of surrogate models in safety-critical contexts—from fusion reactor control to extreme weather response—where uncertainty quantification is imperative but computational budgets preclude ensemble methods or extensive Bayesian inference.
Scope and Limitations. While we focus on spatio-temporal data, our methodology extends to any models producing fixed tensorial outputs with exchangeable calibration and prediction regimes. However, users must carefully validate exchangeability assumptions, particularly for time-series and experimental data where distribution shifts are common. As discussed in Section 4, the marginal nature of coverage guarantees, cell-wise independence assumptions, and input independence in deterministic models represent important limitations that practitioners should consider when applying CP to their specific problems.
Future Directions. Extensions to conditional coverage, incorporation of spatial-temporal correlation structures, and methods for handling systematic exchangeability violations remain important open problems. The integration of CP with recent advances in foundation models for scientific computing presents particularly promising opportunities for scalable, trustworthy uncertainty quantification.
Reproducibility. All code, data generation scripts, and trained models are publicly available at https://github.com/gitvicky/Spatio-Temporal-UQ.
@@ 6 Acknowledgements
The authors would like to thank Anima Anandkumar and Zongyi Li from Caltech for their help with defining neural operators and extending them to complex physics cases. The authors would also like to thank Michael McCabe at the Flatiron Institute for his help in setting up the Multi-Physics Pretrained foundation physics model and extending it to a Fusion-relevant database. This work was performed using resources provided by the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service, provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1), and DiRAC funding from the Science and Technology Facilities Council. This work was supported under project ID a122 as part of the Swiss AI Initiative, through a grant from the ETH Domain and computational resources provided by the Swiss National Supercomputing Centre (CSCS) under the Alps infrastructure. It was also supported by the Excellence Center at Linköping–Lund in Information Technology (ELLIIT). Computations were enabled by the Berzelius resource at the National Supercomputer Centre, provided by the Knut and Alice Wallenberg Foundation. This work has been (part-) funded by the EPSRC Energy Programme [grant number EP/W006839/1]. To obtain further information on the data and models underlying this paper, please contact PublicationsManager@ukaea.uk*.
@@ Appendix A Poisson Equation
The Poisson equation is solved with a finite difference scheme using the py-pde python package (py-pde). Equation 7 is constructed as an initial-value problem, where a scalar uniform field is initialised across the domain and evolved until convergence. A dataset comprising different instances of the 1D Poisson equation is constructed by sampling for different initial values uniformly from within the domain: $u_{\text{init}}\in[0,4)$.
@@
Each MLP is trained to take in the scalar initial field along the 32-point spatial domain to output the final field at the steady state. For the case of STD, the architecture is modified with 1D dropout layers. Each model is trained for up to 1000 epochs using the Adam optimiser adam with a step-decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a quantile loss for the case of CQR and MSE loss in all other cases.
@@ B.1 Physics
As discussed in section 3.2, the dataset is built by solving the one-dimensional Convection Diffusion equation numerically. The physics of the equation, given by the various coefficients, is sampled from a certain range as given in Table 2. Each datapoint, as in each simulation, is generated with different Diffusion coefficients and wave velocities as described in section 3.2. Each simulation is run for 100 time iterations with a $\Delta t=0.0005$ across a spatial domain spanning [0,10], uniformly discretised into 200 spatial units in the x-axis. Once the simulations are run and the dataset is generated, we downsample the temporal discretisation from 100 to 20 by taking every 5th time step. The sampling parameters governing the PDE solutions used for the training are given in Table 2 and that used for the calibration and prediction is given in Table 3.
@@
Diffusion Coefficient $(\alpha)$ $\sin(\frac{x}{\pi}),\sin(\frac{x}{2\pi})]$ Continuous
@@ B.2 Model and Training
We train a U-Net to map the spatio-temporal evolution of the field variable, taking in the first 20 time instances (T_in) to the next 10 time instances (T_out). For the case of the Convection-Diffusion Equation, we don’t deploy an auto-regressive structure but perform a mapping from the initial distribution to the later distribution. The U-Net architecture can be found in Table 4. For the case of STD, the architecture is modified with 1D dropout layers following each encoder and decoder of the U-Net. Though the values governing the evolution of Convection-Diffusion are relatively small, for better representation, we normalise the value with a linear range scaling, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser adam with a step-decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a quantile loss for the case of CQR and an MSE loss in all other cases.
@@
Diffusion Coefficient $(\alpha)$ $\sin(\frac{x}{2\pi}),\sin(\frac{x}{4\pi})]$ Continuous
@@ C.1 Physics
The solution for the wave equation is obtained by deploying a spectral solver that uses a leapfrog method for time discretisation and a Chebyshev spectral method on a tensor product grid for spatial discretisation (GOPAKUMAR2023100464). The dataset is built by performing a Latin hypercube scan across the defined domain for the parameters $\alpha,\beta,\gamma$, which accounts for the amplitude and the location of the Gaussian peak, sampled differently for each simulation. We generate 2,500 simulation points, each one with its own initial condition and use 500 for training, 1,000 each for calibration and procuring the prediction sets. We train a 2D U-Net and an FNO to learn the evolution of wave dynamics.
The physics of the equation, given by the various coefficients, is held constant across the dataset generation throughout, as given in Equation 10. Each data point, as in each simulation, is generated with a different initial condition as described above. The parameters of the initial conditions are sampled from within the domain as given in Table 5. Each simulation is run for 150-time iterations with a $\Delta t=0.00667$ across a spatial domain spanning $1,1]^{2}$, uniformly discretised into 33 spatial units in the x and y axes. Once the simulations are completed and the dataset is generated, we select the first 80 time instances of the evolution of each simulation to be used for training.